Instr. Laura Toma  CPS 130 Questionnaire  May 30th, 2002
What is the pace of the coverage so far?  too slow  just right  too fast 1  2  3  4  5
How much did you get out of the classes?  a little  medium  a lot 1  2  3  4 5
The first homeworks were helpful.  strongly  strongly disagree  agree 1  2  3  4  5

How much time did you spend outside class on homeworks and reading?

a little  medium  a lot
1  2  3  4 5

What were the most important things you learned? What were the least clear/most poorly presented topics? What were the strengths of the classes? What were the weaknesses of the classes? Is there a topic you feel needs to be covered more thoroughly? Feel free to list more than one. What one thing can the instructor of the course do to help improve the class?